Hi...
Could you please help me to solve this problem...
Solve 3 = logx216
I am so confused.
Thank you



Answer by LinnW(1048)

3 = logx(216)
logx(216) means the exponent on x that results in 216.
So we are trying to find x^3 = 216, so x = 6
3 = log6(216) since 6^3 = 216